# The Race between Technology and Woman:

### v8

@@ 7 Conclusion
We developed a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. We applied this framework to developed economies consisting of multiple sectors, where two types of capital (ICT and non-ICT capital) and four types of labor (male skilled, female skilled, male unskilled, and female unskilled labor) are used to produce goods and services.
Our results show that the expansion of ICT equipment results in an increase in the skill wage gap and a decline in the gender wage gap, while the increased supply of male and female skilled labor prevents an increase in the skill wage gap and a decline in the gender wage gap. The observed patterns of changes in the four wage gaps (skilled gender wage gap, unskilled gender wage gap, male skill wage gap, and female skill wage gap) can be explained in terms of the race between progress in ICT and advances in educational attainment and female employment. The unskilled gender wage gap is more likely to decline than the skilled gender wage gap, while the male skill wage gap is more likely to increase than the female skill wage gap. The reason for this is not because the demand for female skilled labor is insufficient but because the supply of female skilled labor becomes abundant. Our results further show that the effects of technological change embodied in ICT equipment are sufficiently large to account for the narrowing of the unskilled gender wage gap, the widening of the male skill wage gap, and about half of the decline in the income share of male unskilled labor, which accounts for most of the decline in the labor share.
@@
Appendix A Appendix
A.1 Proofs of propositions
Proof of Proposition 1
Using the facts that $\partial\ln\mathcal{F}(\ell_{1},...,\ell_{F})/\partial\ln\ell_{f}=\Lambda_{\ell_{f}}$ and $\partial\ln\mathcal{C}(w_{1},...,w_{F},y)/\partial\ln w_{f}=\Lambda_{\ell_{f}}$, the aggregate elasticities of substitution can be rewritten as
@@ A.1 Proofs of propositions
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=$ $\frac{\partial\ln\Lambda_{\ell_{f}}}{\partial\ln w_{g}}-\frac{\partial\ln\Lambda_{g}}{\partial\ln w_{g}}+1.$ (A2)
@@
where $\Psi_{UL}^{w}$, $\Psi_{UR}^{w}$, $\Psi_{LL}^{w}$, and $\Psi_{LR}^{w}$ are $F_{\ell}\times F_{\ell}$, $F_{\ell}\times F_{k}$, $F_{k}\times F_{\ell}$, and $F_{k}\times F_{k}$ matrices, respectively. The equation can be further rewritten as
@@
×$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell}-1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
@@
×$\times\left[\begin{array}[]{ccc}\Phi_{\left(1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Phi_{\left(F_{\ell}-1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(F_{\ell}-1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
@@
×$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell}-1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
@@
×$\times\left[\begin{array}[]{ccc}\Phi_{\left(1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Phi_{\left(F_{\ell}-1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(F_{\ell}-1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
@@
The log-linearization of the profit maximizing condition, $p_{n}=\mathcal{\mathcal{\widetilde{C}}}_{n}(w_{1n},\ldots,w_{Fn})/A_{n}$, yields
@@
The log-linearization of the aggregate price, $p=(\sum_{n=1}^{N}\theta_{cn}^{1/(1-\eta)}p_{n}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$, yields $\sum_{n=1}^{N}\zeta_{n}d\ln(p_{n}/p)=0$. Substituting equation (A21) into this yields
@@
$d\ln\Lambda_{\ell_{f}n}=\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\Pi_{UL,fg\left(n,m\right)}^{w}d\ln\left(\frac{w_{gm}}{p}\right)+\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\Pi_{UR,fg\left(n,m\right)}^{w}d\ln\left(\frac{r_{g}}{p}\right)+\sum_{m=1}^{N}\Theta_{f\left(n,m\right)}^{w}d\ln A_{m}\quad\text{for }f=1,\ldots,F_{\ell}.$
@@
$d\ln\Lambda_{\ell_{f}n}=\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\Pi_{UL,fg\left(n,m\right)}^{w}d\ln\left(\frac{w_{gm}}{p}\right)-\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\Pi_{UR,fg\left(n,m\right)}^{w}d\ln q_{g}+\sum_{m=1}^{N}\Theta_{f\left(n,m\right)}^{w}d\ln A_{m}\quad\text{for }f=1,\ldots,F_{\ell}.$ (A27)
@@
×$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{11}k_{1N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{1N}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{11}k_{F_{k}N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{F_{k}N}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell},N-1}k_{1N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{1N}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell},N-1}k_{F_{k}N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{F_{k}N}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
@@
×$\times\left[\begin{array}[]{ccc}\Theta_{1\left(1,1\right)}^{w}-\Theta_{F_{\ell}\left(N,1\right)}^{w}&\cdots&\Theta_{1\left(1,N\right)}^{w}-\Theta_{F_{\ell}\left(N,N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Theta_{F_{\ell}\left(N-1,1\right)}^{w}-\Theta_{F_{\ell}\left(N,1\right)}^{w}&\cdots&\Theta_{F_{\ell}\left(N-1,N\right)}^{w}-\Theta_{F_{\ell}\left(N,N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
@@
Substituting equation (A29) into this and setting $d\ln\ell_{fn}=0$ yields equation (25).
@@
Substituting equation (A29) into this yields equation (26), where
@@
A.2 Specification issues

### v9

@@ 7 Conclusion
We developed a unified framework to evaluate the effects of capital-embodied technological change on, as well as the contributions of factor inputs to, the relative wages and income shares of different types of labor. We applied this framework to advanced economies consisting of multiple sectors, where two types of capital (ICCT and non-ICCT capital) and four types of labor (male skilled, female skilled, male unskilled, and female unskilled labor) are used to produce goods and services.
Our results show that the expansion of ICCT equipment results in an increase in the skill wage gap and a decline in the gender wage gap, while the increased supply of male and female skilled labor prevents an increase in the skill wage gap and a decline in the gender wage gap. The observed patterns of changes in the four wage gaps (skilled gender wage gap, unskilled gender wage gap, male skill wage gap, and female skill wage gap) can be explained in terms of the race between progress in ICCT and advances in educational attainment and female employment. The unskilled gender wage gap is more likely to decline than the skilled gender wage gap, while the male skill wage gap is more likely to increase than the female skill wage gap. The reason for this is not because the demand for female skilled labor is insufficient but because the supply of female skilled labor becomes abundant. Our results further show that the effects of technological change embodied in ICCT equipment are sufficiently large to account for the narrowing of the unskilled gender wage gap, the widening of the male skill wage gap, and about half of the decline in the income share of male unskilled labor that accounts for most of the decline in the labor share.
@@
Appendix
Appendix A Extension
A.1 Endogenous labor supply
Proposition 4.
$d\ln\left(\frac{w_{f}}{w_{g}}\right)=\sum_{h=1}^{F_{k}}\left(\Xi_{\left(f,h\right)}^{q}-\Xi_{\left(g,h\right)}^{q}\right)d\ln q_{h}+\sum_{n=1}^{N}\left(\Xi_{\left(f,n\right)}^{A}-\Xi_{\left(g,n\right)}^{A}\right)d\ln A_{n}\quad\text{for }f,g=1,\ldots,F_{\ell},$ (A1)
where $\Xi_{\left(f,g\right)}^{q}$ is the $fg$-th element of the $F_{\ell}\times F_{k}$ matrix $\Xi^{q}$, and $\Xi_{\left(f,n\right)}^{A}$ is the $fn$-th element of the $F_{\ell}\times N$ matrix $\Xi^{A}$. The matrices $\Xi^{q}$ and $\Xi^{A}$ are described in Appendix B. Changes in labor shares due to technological change are given by
$d\ln\Lambda_{\ell_{f}}=\sum_{h=1}^{F_{k}}\left\{\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Xi_{\left(f,h\right)}^{q}+\sum_{g\neq f}^{F_{\ell}}\left[-\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}+\sum_{s\neq f,g}^{F}\left(\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{s}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\Xi_{\left(g,h\right)}^{q}\right.\\ \left.-\left[-\left(1-\epsilon_{\ell_{f}k_{h}}^{\mathcal{C}}\right)\Lambda_{k_{h}}+\sum_{s\neq f,h}^{F}\left(\epsilon_{\ell_{f}k_{h}}^{\mathcal{C}}-\epsilon_{\ell_{s}k_{h}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\right\}d\ln q_{h}+\sum_{n=1}^{N}\left\{\Phi_{\left(f,n\right)}^{w}+\sum_{g\neq f}^{F}\left(1-\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}\Xi_{\left(f,n\right)}^{A}\right.\\ \left.+\sum_{g\neq f}^{F_{\ell}}\left[-\left(1-\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{g}}+\sum_{s\neq f,g}^{F}\left(\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}-\epsilon_{\ell_{s}\ell_{g}}^{\mathcal{C}}\right)\Lambda_{\ell_{s}}\right]\Xi_{\left(g,n\right)}^{A}\right\}d\ln A_{n}\quad\text{for }f=1,\ldots,F_{\ell}.$ (A2)
Equations (A1) and (A2) take similar forms to equations (21) and (22), respectively, and nest them as special cases with zero elasticity of labor supply. Most elements of the matrix $\Xi^{q}$ consist of the aggregate elasticities of substitution $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$ only or both the aggregate elasticities of substitution and the aggregate elasticity of labor supply $1/\gamma$, and the others consist of the factor shares of income $\Lambda_{\ell_{f}}$. Equations (A1) and (A2) show that the aggregate elasticities of substitution and the aggregate elasticity of labor supply are sufficient statistics for quantifying the effects of technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.
Corollary 4.
A.2 Imperfect labor mobility
Proposition 5.
$d\ln\left(\frac{\overline{w}_{f}}{\overline{w}_{g}}\right)=\sum_{h=1}^{F_{k}}\sum_{n=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Omega_{f\left(n,h\right)}^{q}-\frac{\Lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\Omega_{g\left(n,h\right)}^{q}\right)d\ln q_{h}\\ +\sum_{n=1}^{N}\sum_{m=1}^{N}\left(\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\Omega_{f\left(n,m\right)}^{A}-\frac{\Lambda_{\ell_{g}n}}{\Lambda_{\ell_{g}}}\Omega_{g\left(n,m\right)}^{A}\right)d\ln A_{m}\quad\text{for }f,g=1,\ldots,F_{\ell},$ (A3)
where $\Omega_{f(n,g)}^{q}$ is the $ng$-th element of the $N\times F_{k}$ matrix $\Omega_{f}^{q}$, and $\Omega_{f(n,m)}^{A}$ is the $nm$-th element of the $N\times N$ matrix $\Omega_{f}^{A}$. The matrices $\Omega_{f}^{q}$ and $\Omega_{f}^{A}$ are described in the proof below. Changes in labor shares due to technological change are given by
$d\ln\Lambda_{\ell_{f}}=\sum_{h=1}^{F_{k}}\sum_{n=1}^{N}\sum_{m=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\left(\sum_{g=1}^{F_{\ell}}\Pi_{UL,fg\left(n,m\right)}^{w}\Omega_{g\left(m,h\right)}^{q}-\Pi_{UR,fh\left(n,m\right)}^{w}\right)d\ln q_{h}\\ +\sum_{n=1}^{N}\sum_{m=1}^{N}\frac{\Lambda_{\ell_{f}n}}{\Lambda_{\ell_{f}}}\left(\sum_{g=1}^{F_{\ell}}\sum_{l=1}^{N}\Pi_{UL,fg\left(n,l\right)}^{w}\Omega_{g\left(l,m\right)}^{A}+\Theta_{f\left(n,m\right)}^{w}\right)d\ln A_{m}\quad\text{for }f=1,\ldots,F_{\ell},$ (A4)
where $\Pi_{UR,fg(n,m)}^{w}$ is the $nm$-th element of the $N\times N$ matrix $\Pi_{UR,fg}^{w}$, $\Pi_{UL,fg(n,m)}^{w}$ is the $nm$-th element of the $N\times N$ matrix $\Pi_{UL,fg}^{w}$, and $\Theta_{f}^{w}$ is the $N\times N$ matrix whose $nm$-th element is given by $\Theta_{f(n,m)}^{w}=-[1-1/(1-\eta)](I_{(n,m)}-\zeta_{m})$. The matrices $\Pi_{UR,fg}^{w}$ and $\Pi_{UL,fg}^{w}$ are described in in the proof below.
Most elements of the matrix $\Omega_{f}^{q}$ consist of the sectoral elasticities of substitution defined as $\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}=\partial\ln(\ell_{fn}/\ell_{gm})/\partial\ln(w_{gm}/w_{fn})$, and the others are the factor shares of income $\Lambda_{\ell_{f}n}$. All elements of the matrices $\Pi_{UR,fg}^{w}$ and $\Pi_{UL,fg}^{w}$ consist of the sectoral elasticities of substitution $\epsilon_{\ell_{fn}\ell_{gm}}^{\mathcal{C}}$ and the factor shares of income $\Lambda_{\ell_{f}n}$. Equations (A3) and (A4) show that the sectoral elasticities of substitution are sufficient statistics for quantifying the effects of capital-embodied technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.
Appendix B Proofs
Proof of Lemma 1
Using the facts that $\partial\ln\mathcal{F}(\ell_{1},...,\ell_{F})/\partial\ln\ell_{f}=\Lambda_{\ell_{f}}$ and $\partial\ln\mathcal{C}(w_{1},...,w_{F},y)$ $\partial\ln w_{f}=\Lambda_{\ell_{f}}$, the aggregate elasticities of substitution can be rewritten as
@@ Appendix B Proofs
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=$ $\frac{\partial\ln\Lambda_{\ell_{f}}}{\partial\ln w_{g}}-\frac{\partial\ln\Lambda_{\ell_{g}}}{\partial\ln w_{g}}+1.$ (A6)
@@
where $\Psi_{UL}^{w}$, $\Psi_{UR}^{w}$, $\Psi_{LL}^{w}$, and $\Psi_{LR}^{w}$ are $_{\ell}\times F_{\ell}, F×F$$_{\ell}\times F_{k}$, $F_{k}\times F_{\ell}$, and $F_{k}\times F_{k}$ matrices, respectively. The equation can be further rewritten as
@@
$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell}-1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
@@
$\times\left[\begin{array}[]{ccc}\Phi_{\left(1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Phi_{\left(F_{\ell}-1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(F_{\ell}-1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
@@
$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell}-1}k_{1}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{1}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell}-1}k_{F_{k}}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}}k_{F_{k}}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
@@
$\times\left[\begin{array}[]{ccc}\Phi_{\left(1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Phi_{\left(F_{\ell}-1,1\right)}^{w}-\Phi_{\left(F_{\ell},1\right)}^{w}&\cdots&\Phi_{\left(F_{\ell}-1,N\right)}^{w}-\Phi_{\left(F_{\ell},N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
Equation (A23) immediately implies equation (A1). Substituting equation (A23) into equation (A18) yields equation (A2).
@@
The log-linearization of the profit maximizing condition, $p_{n}=\mathcal{\mathcal{\widetilde{C}}}_{n}(w_{1n},\ldots,w_{Fn})$ $A_{n}$, yields
@@
The log-linearization of the aggregate price, $p=(\sum_{n=1}^{N}\theta_{cn}^{1/(1-\eta)}p_{n}^{-\eta/(1-\eta)})^{-(1-\eta)/\eta}$, yields $\sum_{n=1}^{N}\zeta_{n}d\ln(p_{n}/p)$ =0=0. Substituting equation (A25) into this yields
@@
$d\ln\Lambda_{\ell_{f}n}=\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\Pi_{UL,fg\left(n,m\right)}^{w}d\ln\left(\frac{w_{gm}}{p}\right)+\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\Pi_{UR,fg\left(n,m\right)}^{w}d\ln\left(\frac{r_{g}}{p}\right)+\sum_{m=1}^{N}\Theta_{f\left(n,m\right)}^{w}d\ln A_{m}\\ \text{for }f=1,\ldots,F_{\ell}.$
@@
$d\ln\Lambda_{\ell_{f}n}=\sum_{g=1}^{F_{\ell}}\sum_{m=1}^{N}\Pi_{UL,fg\left(n,m\right)}^{w}d\ln\left(\frac{w_{gm}}{p}\right)-\sum_{g=1}^{F_{k}}\sum_{m=1}^{N}\Pi_{UR,fg\left(n,m\right)}^{w}d\ln q_{g}+\sum_{m=1}^{N}\Theta_{f\left(n,m\right)}^{w}d\ln A_{m}\\ \text{for }f=1,\ldots,F_{\ell}.$ (A31)
@@
$\times\left[\begin{array}[]{ccc}\epsilon_{\ell_{11}k_{1N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{1N}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{11}k_{F_{k}N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{F_{k}N}}^{\mathcal{C}}\\ \vdots&\vdots&\vdots\\ \epsilon_{\ell_{F_{\ell},N-1}k_{1N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{1N}}^{\mathcal{C}}&\cdots&\epsilon_{\ell_{F_{\ell},N-1}k_{F_{k}N}}^{\mathcal{C}}-\epsilon_{\ell_{F_{\ell}N}k_{F_{k}N}}^{\mathcal{C}}\\ \Lambda_{k_{1}}&\cdots&\Lambda_{k_{F_{k}}}\end{array}\right],$
@@
$\times\left[\begin{array}[]{ccc}\Theta_{1\left(1,1\right)}^{w}-\Theta_{F_{\ell}\left(N,1\right)}^{w}&\cdots&\Theta_{1\left(1,N\right)}^{w}-\Theta_{F_{\ell}\left(N,N\right)}^{w}\\ \vdots&\vdots&\vdots\\ \Theta_{F_{\ell}\left(N-1,1\right)}^{w}-\Theta_{F_{\ell}\left(N,1\right)}^{w}&\cdots&\Theta_{F_{\ell}\left(N-1,N\right)}^{w}-\Theta_{F_{\ell}\left(N,N\right)}^{w}\\ \zeta_{1}&\cdots&\zeta_{N}\end{array}\right].$
@@
Substituting equation (A33) into this and setting $d\ln\ell_{fn}=0$ yields equation (A3).
@@
Substituting equation (A33) into this yields equation (A4), where
@@
Appendix C Additional corollaries
Corollary 5.
Corollary 6.
or
In this case, the labor share can decline with a rise in the capital–labor ratio or the wage–rental price ratio if $\epsilon>1$. Karabarbounis and Neiman (2014) attribute the decline in the labor share to capital-embodied technological change on the basis of findings that the estimated elasticity of substitution exceeds one and that the rental price of capital fell in many countries. Elsby, Hobijn and Şahin (2013) attribute it to other factors on the basis of the observation that the timing of decline in the labor share did not necessarily align with the timing of capital deepening in the United States.
Corollary 7.
Corollary 8.
Suppose that production requires two types of capital (capital equipment, $k_{e}$, and capital structures, $k_{s}$) and two types of labor (skilled labor, $\ell_{h}$, and unskilled labor, $\ell_{u}$), and the aggregate elasticities of substitution satisfy $\epsilon_{\ell_{h}k_{s}}^{\mathcal{F}}=\epsilon_{\ell_{u}k_{s}}^{\mathcal{F}}=1$ and $1/\epsilon_{\ell_{u}\ell_{h}}^{\mathcal{F}}-1/\epsilon_{\ell_{h}\ell_{u}}^{\mathcal{F}}=1/\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}-1/\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}$. Equation (18) becomes the Krusell, Ohanian, Ríos-Rull and Violante (2000) model:
This equation implies that the skill wage gap can increase with a rise in capital equipment if capital equipment is more complementary to skilled labor than to unskilled labor (i.e., $\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}>\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}$). This effect is sufficiently large to account for a rise in the skill premium in OECD countries when ICCT capital is distinguished from non-ICCT capital (Taniguchi and Yamada, 2022).
Appendix D Data Description
D.1 Wages and hours worked
Suppose that there is no need to make an adjustment to wages and hours worked. Let checks denote unadjusted values. The wages for labor of type $f$ in sector $n$, country $j$, and year $t$ could be calculated as $\check{w}_{f,njt}=\sum_{a_{f}}\rho_{f,njt}^{a_{f}}\check{w}_{f,njt}^{a_{f}}$, where $\rho_{f,njt}^{a_{f}}$ is the share of total hours worked by group $a_{f}$ (i.e., $\rho_{f,njt}^{a_{f}}=\check{\ell}_{f,njt}^{a_{f}}/\sum_{a_{f}}\check{\ell}_{f,njt}^{a_{f}}$). The hours worked by labor of type $f$ in sector $n$, country $j$, and year $t$ could be calculated as $\check{\ell}_{f,njt}=\sum_{a_{f}}\check{\ell}_{f,njt}^{a_{f}}$. We adjust for changes in the age and education composition of the labor force by holding the share of each group constant when we calculate wages and by using the time-invariant efficiency units as weights when we calculate hours worked. Let $T_{j}$ denote the number of years observed for country $j$. The composition-adjusted wages for labor of type $f$ in sector $n$, country $j$, and year $t$ can be calculated as $w_{f,njt}=\sum_{a_{f}}\overline{\rho}_{f,nj}^{a_{f}}\check{w}_{f,njt}^{a_{f}}$, where $\overline{\rho}_{f,nj}^{a_{f}}$ is the sector- and country-specific mean of $\rho_{f,njt}^{a_{f}}$ (i.e., $\overline{\rho}_{f,nj}^{a_{f}}=\sum_{t=1}^{T_{j}}\rho_{f,njt}^{a_{f}}/T_{j}$). The composition-adjusted hours worked by labor of type $f$ in sector $n$, country $j$, and year $t$ can be calculated as $\ell_{f,njt}=\sum_{a_{f}}(\overline{w}_{f,nj}^{a_{f}}/\overline{w}_{f,nj}^{a_{f^{\prime}}})\check{\ell}_{f,njt}^{a_{f}}$, where the efficiency unit is measured by the sector- and country-specific mean of $\check{w}_{f,njt}^{a_{f}}$ (i.e., $\overline{w}_{f,nj}^{a_{f}}=\sum_{t=1}^{T_{j}}\check{w}_{f,njt}^{a_{f}}/T_{j}$) and normalized by $\overline{w}_{f,nj}^{a_{f^{\prime}}}$. We choose middle-aged high-skilled labor as the base group for skilled labor and middle-aged medium-skilled labor as the base group for unskilled labor. Our results do not depend on the choice of the base group.
D.2 Rental price of capital
where the interest rate is the internal rate of return:
and the interest rate as
where cpicpi is the consumer price index. The data can be obtained from OECD.Stat.
D.3 Non-ICCT equipment
The trends in the rental prices of capital differ significantly between ICCT equipment and non-ICCT structures but do not differ significantly between non-ICCT equipment and non-ICCT structures (Figure A1a). The rental price of ICCT equipment fell dramatically, but that of non-ICCT equipment and structures remained almost unchanged. Meanwhile, the quantities of ICCT equipment, non-ICCT equipment, and non-ICCT structures increased. However, the rate of increase in ICCT equipment is far greater than that in non-ICCT equipment and structures (Figure A1b). The rate of increase in non-ICCT equipment is almost the same as that in non-ICCT structures.
Figure A1: ICCT equipment, non-ICCT equipment, and non-ICCT structures
(a) Investment prices
(b) Quantities
Notes: The investment prices of ICCT equipment, non-ICCT equipment, and non-ICCT structures are denoted by $p_{i}$, $p_{oe}$, and $p_{os}$, respectively, and the output price is denoted by $p$. The quantities of ICCT equipment, non-ICCT equipment, and non-ICCT structures are denoted by $k_{i}$, $k_{oe}$, and $k_{os}$, respectively. All the series are logarithmically transformed and normalized to zero in the year 1980.
Appendix E Specification Issues